OpenStudy (aundiewear):

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.) sin θ − 1 = cos θ

OpenStudy (kesumonu):

\[\theta= n \pi+(-1)^{n}(\pi/4)+\pi/4\]